Determine the volume of a 0.119 M AlPO4 solution that contains 21.3 grams of AlPO4
Answer #1

Number of moles of AlPO4 = mass of AlPO4 / molar mass of AlPO4

Number of moles of AlPO4 = 21.3 g / 121.9529 g/mol = 0.175 mole

molarity of AlPO4 = number of moles of AlPO4 / volume of solution in L

0.119 = 0.175 / volume of solution in L

volume of solution in L = 0.175 / 0.119 = 1.47 L

1 L = 1000 mL

1.47 L = 1470 mL

Therefore, the volume of AlPO4 = 1470 mL